# Joint Demonstration and Preference Learning Improves Policy Alignment with Human Feedback

### v2

@@ 4 Why AIHF can outperform two-stage alignment approaches?
where $\rho^{*}_{-(i,j)}(r):=\Big{(}1-\sum_{k\in A\backslash\{i,j\}}\pi^{*}_{k}(r)\Big{)}^{-1}$ is the expected number of times an action other than $\tau_{i}$ or $\tau_{j}$ is selected when sampling $\pi^{*}(r)$ infinitely many times. The second-order derivatives are given by
$\frac{\partial^{2}\ell_{\mathcal{P}}(r)}{\partial r_{i}^{2}}=-\frac{1}{\beta^{2}}\sum_{j\neq i}\frac{\pi^{*}_{i}(r)}{\pi^{*}_{i}(r)+\pi^{*}_{j}(r)}\frac{\pi^{*}_{j}(r)}{\pi^{*}_{i}(r)+\pi^{*}_{j}(r)}\frac{|\mathcal{P}_{i,j}|}{|\mathcal{P}|}$ $~{}~{}\frac{\partial^{2}_{r_{i}}\ell_{\mathcal{P}}(r)}{\partial r_{j}\partial r_{i}}=\frac{1}{\beta^{2}}\frac{\pi^{*}_{i}(r)}{\pi^{*}_{i}(r)+\pi^{*}_{j}(r)}\frac{\pi^{*}_{j}(r)}{\pi^{*}_{i}(r)+\pi^{*}_{j}(r)}\frac{|\mathcal{P}_{i,j}|}{|\mathcal{P}|}.$
Note that compared with the previous case with only demonstration data, here the first-order solution does not imply global optimality since the Hessian is negative semi-definite.
Next, by using the above analysis, let us understand the policies that can be obtained by different alignment algorithms.
SFT Estimator. In this case, it is easy to observe that the optimal policy is given by
$\pi^{\rm SFT}_{i}=\frac{\#\tau_{i}~{}\mbox{in}~{}\mathcal{D}}{|\mathcal{D}|}={\pi^{*}(\widehat{r}_{\mathcal{D}})},~{}\forall~{}i.$ (21)
RLHF Policy. The RLHF policy is defined as follows:
where $\widehat{r}_{\mathcal{P}}$ is the estimator obtained from preferences (i.e. solution to (20)), $\pi^{\rm SFT}$ is a prior distribution obtained from the SFT model trained with demonstration dataset $\mathcal{D}$, and $\Delta^{N}$ is the probability simplex. It follows that
$\pi^{\rm RLHF}(\tau_{i})$ $=\frac{\pi^{\rm SFT}(\tau_{i})\exp\frac{1}{\beta}\widehat{r}_{\mathcal{P}}(\tau_{i})}{\sum_{j=1}^{n}\pi^{\rm SFT}(\tau_{j})\exp\frac{1}{\beta}\widehat{r}_{\mathcal{P}}(\tau_{j})}$
$={\frac{\exp\frac{1}{\beta}\widehat{r}_{\mathcal{D}}(\tau_{i})\exp\frac{1}{\beta}\widehat{r}_{\mathcal{P}}(\tau_{i})}{\sum_{j=1}^{n}\exp\frac{1}{\beta}\widehat{r}_{\mathcal{D}}(\tau_{j})\exp{\frac{1}{\beta}\widehat{r}_{\mathcal{P}}}(\tau_{j})}}$
$=\pi^{*}_{i}\Big{(}\widehat{r}_{\mathcal{D}}+\widehat{r}_{\mathcal{P}}\Big{)}$ (22)
That is, the RLHF policy can be seen as the softmax policy for the sum of rewards estimator.
AIHF estimator: In AIHF, assume that the lower-level problem uses (10), and the prior $\pi^{0}$ is uniform, then the optimal $\pi^{\rm AIHF}$ follows the same expression as in (15). One can then plug in such a solution to its loss function in (8a), and obtain the following reward learning problem:
$\widehat{r}^{\rm AIHF}=\arg\max_{r}\quad\ell_{\mathcal{D}+\mathcal{P}}(r):=|\mathcal{D}|L_{1}(r)+|\mathcal{P}|L_{2}(r)$ (23)
where $L_{1}(r):=\mathbb{E}_{\tau_{i}\sim\mathcal{D}}[\log\pi^{*}_{i}(r)]$ and $L_{2}(r):=\mathbb{E}_{(\tau_{j}\prec\tau_{i})\sim\mathcal{P}}[\log\frac{\pi^{*}_{i}(r)}{\pi^{*}_{i}(r)+\pi^{*}_{j}(r)}]$.
Using (16) and (19), we obtain:
$\frac{\partial\ell_{\mathcal{D}+\mathcal{P}}(r)}{\partial r_{i}}$ =#$=\frac{\#\{\tau_{i}~{}\mbox{in}~{}\mathcal{D}\}}{|\mathcal{D}|}|\mathcal{D}|+\sum_{j\neq i}\#\{\tau_{j}\prec\tau_{i}~{}\mbox{in}~{}\mathcal{P}\}-\pi^{*}_{i}|\mathcal{D}|-\sum_{j\neq i}\frac{\pi^{*}_{i}(r)}{\pi^{*}_{i}(r)+\pi^{*}_{j}(r)}|\mathcal{P}_{i,j}|.$
Hence, the first-order condition is:
#$\#\{\tau_{i}~{}\mbox{in}~{}\mathcal{D}\}+\sum_{j\neq i}\#\{\tau_{j}\prec\tau_{i}~{}\mbox{in}~{}\mathcal{P}\}$ $=\pi^{*}_{i}(r)(|\mathcal{D}|+\sum_{j\neq i}|\mathcal{P}_{i,j}|\frac{1}{\pi^{*}_{i}(r)+\pi^{*}_{j}(r)})$
Or equivalently,
$\pi^{*}_{i}(r)$ =#$=\frac{\#\{a_{i}~{}\mbox{in}~{}\mathcal{D}\}+\sum_{j\neq i}\#\{a_{j}\prec a_{i}~{}\mbox{in}~{}\mathcal{P}\}}{|\mathcal{D}|+\sum_{j\neq i}|\mathcal{P}_{i,j}|\rho^{*}_{-(i,j)}(r)}$ (24)
Example 1: With $\beta=1$ and only two actions $\tau_{1}$ and $\tau_{2}$. Since $\rho^{*}(r)_{-(1,2)}=1$, it follows that:
$\pi^{AIHF}_{1}:=\pi^{*}_{1}(\widehat{r}^{AIHF})$ =#$=\frac{\#\{\tau_{1}~{}\mbox{in}~{}\mathcal{D}\}+\#\{\tau_{2}\prec\tau_{1}~{}\mbox{in}~{}\mathcal{P}\}}{|\mathcal{D}|+|\mathcal{P}|}$
$=\frac{|\mathcal{D}|}{|\mathcal{D}|+|\mathcal{P}|}\pi^{*}_{1}(\widehat{r}_{\mathcal{D}})+\frac{|\mathcal{P}|}{|\mathcal{D}|+|\mathcal{P}|}\pi^{*}_{1}(\widehat{r}_{\mathcal{P}})$
That is the policy estimated obtained by using demonstrations and preferences is a weighted average of the polices estimated by using only demonstrations and only preferences. Abusing notation let $\pi^{*}_{1}:=\pi^{*}_{1}(r^{*})$ where $r^{*}$ is the ground-truth reward. It follows that $\mbox{Var}(\pi^{*}_{1}(\widehat{r}_{\mathcal{D}}))=\frac{\pi^{*}_{1}(1-\pi^{*}_{1})}{|\mathcal{D}|}$, $\mbox{Var}(\pi^{*}_{1}(\widehat{r}_{\mathcal{P}}))=\frac{\pi^{*}_{1}(1-\pi^{*}_{1})}{|\mathcal{P}|}$ and
$\mbox{Var}(\pi^{\rm AIHF}_{1})=\frac{\pi^{*}_{1}(1-\pi^{*}_{1})}{|\mathcal{D}|+|\mathcal{P}|}$
This shows the AIHF policy estimator has less variance than either one of the estimators obtained from demonstrations only or preferences only.
To further illustrate, suppose $r(\tau_{1})=r(\tau_{2})$ and we have the following datasets:
With the given data, $\pi^{\rm SFT}_{1}=\pi^{*}_{1}(\widehat{r}_{\mathcal{P}})=\frac{\#\{\tau_{1}~{}\mbox{in}~{}\mathcal{D}\}}{|\mathcal{D}|}=\frac{50}{100}$ and the solution to (20) yields
$\pi^{*}_{1}(\widehat{r}_{\mathcal{P}})=\frac{\exp\widehat{r}_{\mathcal{P}}(\tau_{1})}{\exp\widehat{r}_{\mathcal{P}}(\tau_{1})+\exp\widehat{r}_{\mathcal{P}}(\tau_{2})}=\frac{6}{10}$
$=\frac{\exp\widehat{r}_{\mathcal{P}}(\tau_{1})}{\exp\widehat{r}_{\mathcal{P}}(\tau_{1})+\exp\widehat{r}_{\mathcal{P}}(\tau_{2})}=\frac{6}{10}$
In this example, the RLHF policy estimator is the farthest from ground-truth, because it does not correctly use the information provided by the demonstration data which in this case happens by chance to be correct $\pi^{\rm SFT}_{1}=\pi^{*}_{1}(\widehat{r}_{\mathcal{P}})=\frac{1}{2}$.
Consider now the datasets:
{#$\{\#\tau_{1}~{}\mbox{in}~{}\mathcal{D}\}=6,\quad\{\#\tau_{2}~{}\mbox{in}~{}\mathcal{D}\}=4,\quad\{\#\tau_{1}\succ\tau_{2}~{}\mbox{in}~{}\mathcal{P}\}=\{\#\tau_{2}\succ\tau_{1}~{}\mbox{in}~{}\mathcal{P}\}=50$ (25)
In this case, $\pi^{\rm SFT}_{1}=\pi^{*}_{1}(\widehat{r}_{\mathcal{P}})=\frac{6}{10}$ and the solution to (20) yields
$\pi^{*}_{1}(\widehat{r}_{\mathcal{P}})=\frac{\exp\widehat{r}_{\mathcal{P}}(\tau_{1})}{\exp\widehat{r}_{\mathcal{P}}(\tau_{1})+\exp\widehat{r}_{\mathcal{P}}(\tau_{2})}=\frac{50}{100}$
It follows from (22) that:
$=\frac{\pi_{1}^{SFT}}{\pi_{1}^{SFT}+\pi_{2}^{SFT}}=\frac{6}{10}$
Numerical Example 2. Let us use an illustrative example to show that RLHF method will result in significant data under-utilization when the demonstration coverage is limited. Assume that there are 50 actions, each with a ground-truth reward defined by $r_{i}=\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac{(\frac{i}{50}-\mu)^{2}}{2\sigma^{2}}}$, where $\mu=0.5$ and $\sigma=2$. Assume we can sample demonstration and preference from the ground truth reward distribution: demonstrations are sampled from the multinomial distribution, while preferences are sampled from the BTL model.
In an extreme scenario, let demonstrations cover actions 1 through 45, while preferences have full coverage across all actions. In the subsequent experiment, we initially sample 2000 demonstrations using the multinomial distribution $p_{i}=\frac{\exp r_{i}}{\sum^{j=45}_{j=1}\exp r_{j}}$, and obtain 200 preferences for each preference pair with $p(i\succ j)=\frac{\exp r_{i}}{\exp r_{j}+\exp r_{i}}$. We then calculate the optimal policy by solving the previously mentioned first order conditions, yielding the following result.
From the result shown in Figure 2, we demonstrate that both SFT and RLHF transfer the weight from uncovered action to covered action when demonstration coverage is limited, as indicated by $\pi_{SFT}(a_{uncovered})=0$. Consequently, the weight of covered actions is significantly higher than the ground truth. However, this issue does not occur when jointly optimizing the demonstration and preference in the AIHF method.
@@ 5 Proposed Algorithm for AIHF Training
We are now ready to design algorithms for the proposed AIHF formulation (8). To begin with, first note that (8) takes a hierarchical form, and it belongs to the class of problem named bi-level optimization, first developed in the 70s (Fiacco and McCormick, 1990), and recently found many applications in machine learning (Wang et al., 2021; Liu et al., 2021, 2022). Generically speaking, bi-level problems are not easy to optimize; more specifically, in (8), the upper-level problem (8a) is a function of both the lower-level optimal solution $\pi_{\theta}$ and the true parameter $\theta$. It follows that a (stochastic) first-order algorithm for $L(\theta)$ involves some (potentially non-trivial) implicit gradient computation which often involves computing the Hessian matrix for the lower-level objective function. Fortunately, as we will show shortly, with some special choices of $L_{1}$, $L_{2}$, $L_{3}$, one can design some simple and very efficient algorithms.
Throughout this section, we are searching for a good policy $\pi_{\theta}$ and a reward estimate $r(\cdot,\cdot;\theta)$ to align with human feedback, where the policy $\pi_{\theta}$ is an optimal solution w.r.t. the certain reward estimate $r(\cdot,\cdot;\theta)$ according to the policy optimization problem (3.4). Due to such optimal policy constraint w.r.t. one explicit reward estimate, we design an algorithm to solve such a single-stage, bi-level problem which is different from DPO (Rafailov et al., 2023) that simply optimizes the fixed loss function (13) directly.
@@
Policy Alignment Step. From our earlier discussion, we know that when $L_{3}$ takes the form (10), the optimal policy (for a fixed reward) is given by (5). Of course, one cannot directly compute such an optimal solution due to the fact that both $Q_{\theta}$ and the normalization term are unknown. Therefore one can adopt the standard approaches such as the well-known proximal policy optimization (PPO) (Schulman et al., 2017) algorithm to obtain an approximate optimal policy. It is worth noting that, when considering $T=1$, our discussion leading to (7) indicates the optimal policy takes a much simpler form. In this case, it is possible to consider simpler method than running PPO to obtain the optimal policy. One alternative way is to use a baseline estimated reward value to perform variance reduction Li et al. (2023), thus reducing the computational complexity.
@@
Reward Alignment. In this step, we use a stochastic gradient-type algorithm to optimize $L(\theta)$. Towards this end, first observe that
$\nabla L(\theta)$ $=\alpha\nabla L_{1}(\pi_{\theta})+\nabla L_{2}(\theta).$ (26)
@@
$\nabla{L}^{\rm BTL}_{2}(\theta)$ $=\mathbb{E}_{(\tau_{l}\prec\tau_{w})\sim\pi^{P}}\Big{[}\nabla_{\theta}\log\Big{(}\sigma\big{(}R(\tau_{w};\theta)-R(\tau_{l};\theta)\big{)}\Big{)}\Big{]}.$ (27)
On the contrary, the computation of $\nabla L_{1}(\pi_{\theta})$ is more involved, since $L_{1}$ depends on $\theta$ implicitly through the corresponding optimal policy $\pi_{\theta}$. Fortunately, the following lemma indicates that this gradient has a simple and intuitive form as well.
@@
In practice, a few approximations need to be made to obtain a stochastic gradient of $L_{1}$. First, similarly, as before, the precise expectation cannot be obtained because the ground truth policy $\pi^{\rm E}$ is unknown. Denote an offline demonstration dataset as $\mathcal{D}^{\rm E}:=\{\tau\}$, then one can replace the expectations $\mathbb{E}_{\tau\sim\pi^{E}}$ by $\mathbb{E}_{\tau\sim\mathcal{D}^{\rm E}}$. Second, in the second expectation in (28), the trajectories $\tau^{\prime}$ are sampled from $\pi_{\theta}$, the optimal policy for a fixed reward parameterization by $\theta$. This means that the policy alignment step has to identify the optimal policy $\pi_{\theta}$ first, which, due to limitations such as computational constraints, and non-linear parameterization, is generally not possible. Instead, we propose to sample from the current policy $\pi^{k+1}$ obtained from the previous policy optimization step, where index $k$ represents the iteration counter. The above two steps are summarized in Algorithm 1.
Let us remark on the gradient expression of the proposed algorithm. Here, we explain how to construct the gradient estimator $g^{k}_{1}$ for $\nabla_{\theta}L_{1}(\theta_{k})$ defined in (28) and $g^{k}_{2}$ for $\nabla_{\theta}L_{2}(\theta_{k})$ defined in (27). At iteration $k$, to estimate $\nabla_{\theta}L_{1}(\theta_{k})$, we need to sample $N$ expert trajectories $\tau^{\rm E}$ from the expert demonstration dataset $\mathcal{D}^{\rm E}$ and $N$ agent trajectories $\tau^{\prime}$ from the policy $\pi^{k+1}$. Then the gradient estimator $g_{1}^{k}$ can be constructed as: $g_{1}^{k}:=\frac{1}{N}\sum_{i=1}^{N}\Big{[}\nabla_{\theta}R(\tau_{i};\theta)-\nabla_{\theta}R(\tau_{i}^{\prime};\theta)\Big{]}$. Furthermore, through sampling $M$ pairs of preference $(\tau_{l}\prec\tau_{w})$, we can further construct the gradient estimator of $\nabla_{\theta}L_{2}(\theta_{k})$ as: $g_{2}^{k}=\sum_{i=1}^{M}\nabla_{\theta}\log\Big{(}\sigma\big{(}R(\tau_{i,w};\theta)-R(\tau_{i,l};\theta)\big{)}\Big{)}$. Here, as a remark, although the demonstration data is used, but in the final algorithm, the gradient is always in constrastive form. The difference is that the for the demonstration data, it is compared with its own model generation, while for preference data, two preference data is compared.
Before we close this section, let us remark on the computational complexity of the proposed algorithm. Note that our algorithm is motivated by a class of popular algorithms in bi-level optimization, where the upper-level and lower-level problems are updated alternatingly using stochastic optimization (Hong et al., 2020). Specifically, one only needs to perform one (or a few steps) of policy optimization step (such as PPO) to obtain an improved $\pi^{k+1}$ before moving to the next iteration. As demonstrated in our subsequent numerical results, the computational complexity of our method is comparable or sometimes even lower as compared to their sequential optimization counterparts such as RLHF. In our numerical results, we show that the proposed algorithm AIHF can achieve a comparable performance as RLHF while performing less policy optimization steps.
@@
  Input: Initialize reward parameter $\theta^{0}$ and policy model $\pi^{0}$, the stepsize of reward update $\eta$. Let $\mathcal{P}$, $\mathcal{D^{\rm E}}$ denote the preference and the demonstration data, respectively.   for  Iteration $k=0,1,\ldots,K-1$ do      Policy Alignment: Optimizing $L_{3}$ by RL subroutine, e.g. PPO, to obtain one improved policy $\pi^{k+1}$      Data Sample I: Sample expert trajectories $\tau\sim\mathcal{D^{\rm E}}$ and agent trajectories from $\tau^{\prime}\sim\pi^{k+1}$      Data Sample II: Sample preference pair $(\tau_{l}\prec\tau_{w})\sim\mathcal{P}$      Estimating Gradient: Calculate the estimator $g^{k}:=\alpha g_{1}^{k}+g_{2}^{k}$ of $\nabla_{\theta}L(\theta_{k})=\alpha\nabla_{\theta}L_{1}(\theta_{k})+\nabla_{\theta}L_{2}(\theta_{k})$      Reward Alignment: $\theta^{k+1}:=\theta^{k}+\eta g^{k}$   end for

### v3

@@ 3.4 Why AIHF can outperform two-stage alignment approaches?
where $|\mathcal{P}_{i\succ j}|:=\#\{\tau_{i}\succ\tau_{j}~{}\mbox{in}~{}\mathcal{P}\}$ and $|\mathcal{P}_{i,j}|:=|\mathcal{P}_{i\succ j}|+|\mathcal{P}_{j\succ i}|$ and $\rho_{-(i,j)}(\pi):=\Big{(}1-\sum_{k\in A\backslash\{i,j\}}\pi_{k}\Big{)}^{-1}$ is the expected number of times an action other than $\tau_{i}$ or $\tau_{j}$ is selected when sampling actions from $\pi$ infinitely many times.
RLHF Policy. Based on the above results, in Sec. A.2.2 we show that the RLHF approach has the optimal policy $\pi^{\rm RLHF}(\tau_{i})=\pi^{*}_{i}\Big{(}\widehat{R}_{\mathcal{D}}+\widehat{R}_{\mathcal{P}}\Big{)}$. That is, the RLHF policy can be seen as the softmax policy for the sum of reward estimators obtained from demonstrations and preferences separately.
AIHF Policy. Finally, we also find that the AIHF policy is of the form:
$\pi^{\rm AIHF}(\tau_{i})=\frac{\#\{\tau_{i}~{}\mbox{in}~{}\mathcal{D}\}+\sum_{j\neq i}|\mathcal{P}_{i\succ j}|}{|\mathcal{D}|+\sum_{j\neq i}|\mathcal{P}_{i,j}|\rho_{-(i,j)}\big{(}\pi^{*}(\widehat{R}^{\rm AIHF})\big{)}}.$ (15)
Discussion. Let us summarize our findings. First, $\pi^{\rm RLHF}$, which takes the form of the softmax of the sum of two rewards—one learned during the SFT stage and another during the reward training stage—offers some interesting insights into this popular approach. Second, $\pi^{\rm AIHF}$ is more robust than $\pi^{\rm RLHF}$. To see this, suppose that $|\mathcal{D}|\gg|\mathcal{P}|$, i.e. there is more demonstration than preference data. In this case, the policy estimator in (15) will be largely defined by the demonstration data (which is reasonable) whereas the RLHF policy can be noisy since it (soft) maximizes the sum of two reward estimators: one that is more accurate (i.e. the one based on demonstrations, $\widehat{R}_{\mathcal{D}}$) and one that is less accurate (i.e. the one based on preferences $\widehat{R}_{\mathcal{P}}$). A similar observation can be made when $|\mathcal{D}|\ll|\mathcal{P}|$; see Sec. A.2.2 for more details. Third, $\pi^{\rm AIHF}$ can have less variance as compared with $\pi^{\rm RLHF}$. Indeed, $\pi^{\rm AIHF}$ takes a form of a weighted average of the policies estimated separately with demonstration and preference data, as by using (12) and (14), we can re-write (15) as:
$\pi^{*}_{i}(\widehat{R}^{\rm AIHF})=\frac{|\mathcal{D}|}{|\mathcal{D}|+\sum_{j\neq i}|\mathcal{P}_{i,j}|\rho_{-(i,j)}\big{(}\pi^{*}(\widehat{R}^{\rm AIHF})\big{)}}\pi^{*}_{i}(\widehat{R}_{\mathcal{D}})+\frac{\sum_{j\neq i}|\mathcal{P}_{i,j}|\rho_{-(i,j)}\big{(}\pi^{*}(\widehat{R}_{\mathcal{P}})\big{)}}{|\mathcal{D}|+\sum_{j\neq i}|\mathcal{P}_{i,j}|\rho_{-(i,j)}\big{(}\pi^{*}(\widehat{R}^{\rm AIHF})\big{)}}\pi^{*}_{i}(\widehat{R}_{\mathcal{P}})$
Such averaging entails reduced variance. We also include simple numerical examples in the Appendix A.2.3 to further illustrate this point.
@@ 4 Proposed Algorithm for AIHF Training
We are now ready to design algorithms for the proposed AIHF formulation (5). To begin with, first note that (5) takes a hierarchical form, and it belongs to the class of problem named bi-level optimization, first developed in the 70s (Fiacco & McCormick, 1990), and recently found many applications in machine learning (Wang et al., 2021; Liu et al., 2021, 2022). Generically speaking, bi-level problems are not easy to optimize; more specifically, in (5), the upper-level problem (5a) is a function of both the lower-level optimal solution $\pi_{\theta}$ and the true parameter $\theta$. It follows that a (stochastic) first-order algorithm for $L(\theta)$ involves some (potentially non-trivial) implicit gradient computation which often involves computing the Hessian matrix for the lower-level objective function. Fortunately, as we will show shortly, with some special choices of $L_{1}$, $L_{2}$, and $L_{3}$, one can design some simple and very efficient algorithms.
Before we go to details, we note that throughout this section, we assume that we are searching for a good policy $\pi_{\theta}$ and a reward estimate $r(\cdot,\cdot;\theta)$ to align with human feedback, where the policy $\pi_{\theta}$ is an optimal solution w.r.t. the certain reward estimate $r(\cdot,\cdot;\theta)$ according to the policy optimization problem (5b). Due to such optimal policy constraint w.r.t. one explicit reward estimate, we design an algorithm to solve such a single-stage, bi-level problem which is different from DPO (Rafailov et al., 2023) that simply optimizes the fixed loss function (10) directly.
@@
Policy Alignment Step. One can adopt the standard approaches such as the well-known proximal policy optimization (PPO) (Schulman et al., 2017) algorithm to obtain an approximate optimal policy that solves (7). It is worth noting that, when considering $T=1$, our discussion leading to (42) indicates the optimal policy takes a much simpler form. In this case, it is possible to consider a simpler method than running PPO to obtain the optimal policy. One alternative way is to use a baseline estimated reward value to perform variance reduction (Li et al., 2023), thus reducing the computational complexity.
@@
Reward Alignment. In this step, we use a stochastic gradient-type algorithm to optimize $L(\theta)$. Towards this end, first, observe that
$\nabla L(\theta)$ $=w_{1}\nabla L_{1}(\pi_{\theta})+\nabla L_{2}(\theta).$ (16)
@@
$\nabla{L}^{\rm BTL}_{2}(\theta)$ $=\mathbb{E}_{(\tau_{w}\succ\tau_{l})\sim\pi^{P}}\Big{[}\nabla_{\theta}\log\Big{(}\sigma\big{(}R(\tau_{w};\theta)-R(\tau_{l};\theta)\big{)}\Big{)}\Big{]}.$ (17a)
On the contrary, the computation of $\nabla L_{1}(\pi_{\theta})$ is more involved, since $L_{1}$ depends on $\theta$ implicitly through the corresponding optimal policy $\pi_{\theta}$. Fortunately, the following lemma indicates that this gradient has a simple and intuitive form as well, and the proof can be found in Appendix A.3.2.
@@
In practice, a few approximations need to be made to obtain a stochastic gradient of $L_{1}$. First, similarly, as before, the precise expectation cannot be obtained because the ground truth policy $\pi^{\rm E}$ is unknown. Denote an offline demonstration dataset as $\mathcal{D}^{\rm E}:=\{\tau\}$, then one can replace the expectations $\mathbb{E}_{\tau\sim\pi^{E}}$ by $\mathbb{E}_{\tau\sim\mathcal{D}^{\rm E}}$. Second, in the second expectation in (18), the trajectories $\tau^{\prime}$ are sampled from $\pi_{\theta}$, the optimal policy for a fixed reward parameterization by $\theta$. This means that the policy alignment step has to identify the optimal policy $\pi_{\theta}$ first, which, due to limitations such as computational constraints, and non-linear parameterization, is generally not possible. Instead, we propose to sample from the current policy $\pi^{k+1}$ obtained from the previous policy optimization step, where index $k$ represents the iteration counter. Following the approximation steps mentioned above, we construct a stochastic estimator $g_{k}$ to approximate the exact gradient $\nabla L(\theta_{k})$ in (16) as follows:
$g_{k}:=w_{1}g^{k}_{1}+g^{k}_{2}:=$ $w_{1}\big{(}\nabla_{\theta}R(\tau^{E}_{k},\theta_{k})-\nabla_{\theta}R(\tau^{A}_{k},\theta_{k})\big{)}+\big{(}1-\sigma(R(\tau^{W}_{k},\theta_{k})-R(\tau^{L}_{k},\theta_{k}))\big{)}$
×$\quad{\times}\big{(}\nabla_{\theta}R(\tau^{W}_{k},\theta_{k})-\nabla_{\theta}R(\tau^{L}_{k},\theta_{k})\big{)}.$ (19)
The above two steps are summarized in Algorithm 1. let us remark on the computational complexity of the proposed algorithm. Note that our algorithm is motivated by a class of popular algorithms in bi-level optimization, where the upper-level and lower-level problems are updated alternatingly using stochastic optimization (Hong et al., 2023). We conclude the section by theoretically inspecting the proposed algorithms.
@@
  Input: Initialize reward parameter $\theta^{0}$ and policy model $\pi^{0}$, the stepsize of reward update $\eta$. Let $\mathcal{P}$, $\mathcal{D^{\rm E}}$ denote the preference and the demonstration data, respectively.   for  Iteration $k=0,1,\ldots,K-1$ do      Policy Alignment: Optimizing $L_{3}$ by RL subroutine, e.g. PPO, to obtain one improved policy $\pi^{k+1}$      Data Sample I: Sample an expert trajectory $\tau\sim\mathcal{D^{\rm E}}$ and agent trajectory from $\tau^{\prime}\sim\pi^{k+1}$      Data Sample II: Sample preference pair $(\tau_{w}\succ\tau_{l})\sim\mathcal{P}$      Estimating Gradient: Calculate one gradient estimator $g^{k}:=w_{1}g_{1}^{k}+g_{2}^{k}$ of $\nabla_{\theta}L(\theta)=w_{1}\nabla_{\theta}L_{1}(\theta)+\nabla_{\theta}L_{2}(\theta)$      Reward Alignment: $\theta^{k+1}:=\theta^{k}+\eta g^{k}$   end for
@@
Theorem 4.1
Suppose Assumptions 1 - 2 hold. Selecting stepsize $\alpha:=\frac{\alpha_{0}}{K^{\sigma}}$ for the reward update step (4) where $\alpha_{0}>0$ and $\sigma\in(0,1)$ are some fixed constants, and $K$ is the total number of iterations to be run by the algorithm. Then the following result holds:
$\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\left[\big{\|}\log\pi_{k+1}-\log\pi_{\theta_{k}}\big{\|}_{\infty}\right]=\mathcal{O}(K^{-1})+\mathcal{O}(K^{-\sigma})$ (20a)
$\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\left[\|\nabla L(\theta_{k})\|^{2}\right]=\mathcal{O}(K^{-\sigma})+\mathcal{O}(K^{-1+\sigma})+\mathcal{O}(K^{-1})$ (20b)
where $\|\log\pi_{k+1}-\log\pi_{\theta_{k}}\|_{\infty}:=\max_{s\in\mathcal{S},a\in\mathcal{A}}\big{|}\log\pi_{k+1}(a|s)-\log\pi_{\theta_{k}}(a|s)\big{|}$. In particular, setting $\sigma=1/2$, then both quantities in (20a) and (20b) converge with the rate $\mathcal{O}(K^{-1/2}).$
The above theorem shows that Alg. 1 could converge to the stationary point if we take a large loop number $K$. Note that details and proofs of the result above are delegated to Appendix A.3.